OpenStudy (anonymous):

How do I find the square root of 99 if without having a decimal answer?

OpenStudy (dape):

You can't, it's an irrational number. This is because \(\sqrt{99}=\sqrt{3^2*11}=3\sqrt{11}\). The number 11 is prime, any root of a prime is irrational, as such, the root of 99 also is irrational.